OpenStudy (anonymous):

identify the slope and intercept of the following linear equation y=3/7x-5 A)Slope:3/7; intercept:-5 B)Slope:5; intercept:3/7 C)Slope:3/7; intercept:5 D)Slope:-5; intercept:3/7

OpenStudy (anonymous):

A.

OpenStudy (jhannybean):

First draw out the line 

OpenStudy (jhannybean):

this equation is in the slope- intercept form of \(\ y=mx+b\) where \(\ m\) is the slope.

OpenStudy (jhannybean):

\[\ m= \frac{rise}{run} = \frac{up \ \ 3}{over \ \ 7}\]